OpenStudy (anonymous):

The perimeter of a triangle is 40 inches. The second side exceeds twice the first side by 1 inch, and the third side is 2 inches less than the second side. Find the length of each side of the triangle

OpenStudy (anonymous):

please help anyone?! :)

OpenStudy (compassionate):

I can help if you promise me some chocolate.

OpenStudy (zepp):

Let the first side's length be x; and let's proceed sentence by sentence.

OpenStudy (zepp):

The second side exceeds twice the first side by 1 inch. How would you write this algebraically?

OpenStudy (anonymous):

umm...

OpenStudy (anonymous):

2+x+1?

OpenStudy (anonymous):

im not suree =/

OpenStudy (zepp):

Twice implies multiplying, you are on the right track =D

OpenStudy (zepp):

Keep trying, you'll get it ;)

OpenStudy (anonymous):

2x+x+1? im still not sure :(

OpenStudy (jiteshmeghwal9):

Let the first side be x.then second side=x+1 third side=(x+1)-2 A/q, (x)+(x+1)+((x+1)-2)=40 x+x+1+x-1=40 3x=40 x=40/3

OpenStudy (zepp):

The second side exceeds twice the first side by 1 inch. Twice the first side means -> 2x Exceeds by 1 inch implies that's 1 inch more -> 2x+1 Got it?

OpenStudy (anonymous):

ohhhh i get it

OpenStudy (zepp):

Alright so now we have First side = x Second side = 2x+1; Now, the third side is 2 inches less than the second side. How would you write that algebraically?

OpenStudy (anonymous):

2x+1-2?

OpenStudy (zepp):

Excellent! We can simplify that: \(2x-1\)

OpenStudy (anonymous):

oh ok

OpenStudy (zepp):

Alright, now we got First side = x Second side = 2x+1 Third side = 2x - 1 What's the definition of a perimeter?

OpenStudy (anonymous):

40

OpenStudy (zepp):

I mean, the definition of 'perimeter'

OpenStudy (anonymous):

when you add up all the sides?

OpenStudy (zepp):

That's right, so in our case; Perimeter = 1st side + 2nd side + 3rd side Now you know what to do! :)

OpenStudy (jiteshmeghwal9):

Perimeter means the sum of all the sides . & this case is done with every geometrical shape

OpenStudy (anonymous):

okk i'll remember that

OpenStudy (zepp):

So? What's your equation? :D

OpenStudy (anonymous):

well i got stumped on it..

OpenStudy (anonymous):

no nevermind!

OpenStudy (zepp):

Got it? :D

OpenStudy (anonymous):

i made a slight error thats what messed me up

OpenStudy (anonymous):

is it 10?

OpenStudy (zepp):

x = 10?

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

ahh 8

OpenStudy (zepp):

Let's see, definition of 'perimeter': sum of all sides of a geometrical shape. In our question, it would be the sum of all 3 sides First side = x Second side = 2x+1 Third side = 2x - 1 (x)+(2x+1)+(2x-1) would be our perimeter, which is 40 (x)+(2x+1)+(2x-1) = 40 x + 2x + 2x + 1 + 1 = 40 5x + 2 = 40 5x = 40 - 2 5x = 38 x = 38 \(\div\)5 x = 7.6

OpenStudy (zepp):